MATH 135 – Calculus 1 Derivative Rule Practice October 11, 2013

Directions: Use the sum, constant multiple, product, and quotient rules (only) to compute each of the following derivatives. (If you know the Chain Rule, please do not make use of that since we will only learn that in a few days(!)) Say which rules you are using for each one.

I. Find f′(x) for

f(x) = x^{4}e^{x} − \frac{3}{\sqrt{x}}

II. Find G′(t) for

G(t) = e^{−t}(t^{2} − 4t + 3)

III. Find H′(z) and H′′(z) for

H(z) = \frac{z^{2}}{z^{2} − 4}

IV. Find J′(w) for

J(w) = \frac{w^{4} − 7w^{3} + 2w − 1}{w^{3}}